OpenStudy (anonymous):

The sun roof in Troy’s house is shaped like a rectangle. He increases the length of the sides of the sun roof to four times the existing ones. How will this change affect the perimeter of the sun roof? It will be 8 times the original perimeter. It will be 2 times the original perimeter. It will be 32 times the original perimeter. It will be 4 times the original perimeter.

OpenStudy (anonymous):

i chose D

OpenStudy (anonymous):

Rectangle has 4 sides right?

OpenStudy (anonymous):

yes, a rectangle has 4 sides

OpenStudy (anonymous):

Exactly, so dont let this question fool you, because its actually very easy, just a trick question

OpenStudy (anonymous):

so was my answer right or wrong? I'm pretty sure it was

OpenStudy (mathstudent55):

 The perimeter was 2(a + b). Now it is 8(a + b). 8(a + b) / ( 2 (a + b) ) = 8/2 = 4